# Understanding Self-Regulated Learning Behavior Among High and Low Dropout Risk Students During CS1: Combining Trace Logs, Dropout Prediction and Self-Reports

### v1

@@ Abstract.
The introductory programming course (CS1) at the university level is often perceived as particularly challenging, contributing to high dropout rates among Computer Science students. Identifying when and how students encounter difficulties in this course is critical for providing targeted support. This study explores the behavioral patterns of CS1 students at varying dropout risks using self-regulated learning (SRL) as the theoretical framework. Using learning analytics, we analyzed trace logs and task performance data from a virtual learning environment to map resource usage patterns and used student dropout prediction to distinguish between low and high dropout risk behaviors. Data from 47 consenting students were used to carry out the analysis. Additionally, self-report questionnaires from 29 participants enriched the interpretation of observed patterns. The findings reveal distinct weekly learning strategy types and categorize course behavior. Among low dropout risk students, three learning strategies were identified that different in how students prioritized completing tasks and reading course materials. High dropout risk students exhibited nine different strategies, some representing temporary unsuccessful strategies that can be recovered from, while others indicating behaviors of students on the verge of dropping out. This study highlights the value of combining student behavior profiling with predictive learning analytics to explain dropout predictions and devise targeted interventions. Practical findings of the study can in turn be used to help teachers, teaching assistants and other practitioners to better recognize and address students at the verge of dropping out.
@@ 1. Introduction
Identifying struggling students and those at risk of dropping out is crucial to ensuring these interventions are effective. Academic performance prediction and student dropout detection have been extensively researched (Hellas et al., 2018; Prenkaj et al., 2020). Many studies apply learning analytics techniques to various student data, including demographics (e.g., Naseem et al., 2019), basic activity indicators (e.g., Waheed et al., 2023), outcomes of student assignments (e.g., Zhidkikh et al., 2024) and self-reports (e.g., Tempelaar et al., 2018). Notably, Moreno-Marcos et al. (2020) evaluated an effective prediction approach using students’ self-regulation behavior data collected from self-reports and trace logs. Although most student dropout prediction methods demonstrate high accuracy, easy explainability of the predictions remains a challenge (Zhidkikh et al., 2024).
@@
To increase the effectiveness of the predictions and interventions, it is therefore essential to understand how students study and how they utilize available learning resources. Recent research has increasingly focused on identifying student learning behaviors, emphasizing the self-regulation, metacognition, and motivation aspects (Lorås et al., 2022). For instance, common activity indicators can be used to model students’ learning approaches and to identify student behavior profiles (e.g., Kim et al., 2020; Sharma et al., 2015). Similarly, trace logs from virtual learning environments can reveal patterns in students’ learning resource usage and detect common learning strategies used in a classroom (e.g., Jovanović et al., 2017; Zhidkikh et al., 2023). Additionally, self-reports, such as questionnaires, learning journals, and interviews may provide insights into students’ use of learning strategies and cognitive processes (e.g., Loksa and Ko, 2016; Loksa et al., 2020; Silva et al., 2024). However, while these methods help map learning behaviors, they usually do not directly link behavior to early dropout (Zhidkikh et al., 2023). At most, some studies have compared behaviors of high and low-performing students, often in a post-course setting. For instance, Liao et al. (2019) interviewed 19 CS1 students who took the final CS1 exam and identified general differences in learning paths between higher-performing and lower-performing students.
@@
Motivated by the examples above, this paper aims to deepen the understanding of learning behaviors among high and low-risk students during CS1. The present paper complements prior key studies (Liao et al., 2019; Moreno-Marcos et al., 2020). Specifically, we map students’ learning paths, extending previous work (cf. Liao et al., 2019) by tracking weekly learning behaviors using trace logs collected during the course. Further, we employ dropout prediction to identify at-risk students for every course week (cf. Moreno-Marcos et al., 2020) and use this information to discern weekly learning behavior by the associated dropout risk. To this end, the paper presents a learning analytics-based method to link learning resource usage patterns with dropout predictions, combining tools developed in prior studies (Zhidkikh et al., 2023, 2024). A post-course self-report questionnaire was also administered to support the interpretation of the behavioral patterns and to evaluate the potential of the learning analytics used. Both qualitative and quantitative data were inductively explored to gain insights into students’ behaviors at different dropout risk levels. We use self-regulated learning as the foundational theoretical framework widely applied in CSEd research to understand the learning process (Prather et al., 2020). The study ultimately seeks to answer the following research questions:
@@ 2.1. Self-regulated learning and learning strategies
Self-regulation generally refers to the ability to monitor and manage one’s effort, time, motivation and behavior when learning or problem-solving (Pintrich, 2000; Schunk, 2005). While metacognition describes student’s knowledge to self-evaluate learning outcomes and approaches, self-regulation describes the ability to control them (Prather et al., 2020). Therefore, SRL is a strategic learning process in which a student self-regulates by actively setting goals, engaging with the learning resources using various strategies, monitoring progress and refining their approach based on the outcomes and external learning context (Zimmerman, 2008; Winne and Hadwin, 1998; Winne and Marzouk, 2019). Although multiple models detailing the SRL process exist (Panadero, 2017; Loksa et al., 2022), SRL is often described as a cyclical process consisting of forethought and goal-setting, enacting on the goals and self-controlling, and self-reflection on the outcomes and the process (Zimmerman and Moylan, 2009). While the general SRL process can be modelled, the specific ways a student regulates their behavior can depend highly on the context, such as the subject, student’s prior experiences, the task at hand and even other students (Ben-Eliyahu and Bernacki, 2015).
@@
Various measures of SRL have been applied in the academic contexts. Common measures include self-report tools that assess student’s aptitudes for self-regulation (e.g., questionnaires, think-aloud protocols, learning journals), and event measures that map SRL processes via observable, externally collected data (Roth et al., 2016). While some researchers have noted issues related to the reliability of self-reports compared to event-based measures (e.g., Choi et al., 2023; Zimmerman, 2008), a growing number of studies has argued for the combined use of both measures (e.g., Fan et al., 2023; Lim et al., 2021; Tempelaar et al., 2020). As Roth et al. (2016) summarize, the triangulation of new and established measures enables researchers to gain new insights into SRL.
@@
A common aspect used to evaluate and study SRL is the regulation of behavior through learning strategies. A learning strategy is generally a set of planned actions, such as rehearsing, seeking assistance and reviewing course materials, aimed to acquire information and skills relevant to the goal or task at hand (Zimmerman and Martinez-Pons, 1986). For event-based measures, a more fine-grained description of learning strategies, as proposed by Winne (2001) and elaborated by Winne and Marzouk (2019), can be used. In their model, student behavior can be described through simple, atomic actions, such as reading a text section or writing a note. A sequence of actions aimed at completing a specific sub-goal or a task comprises a learning tactic. For example, a student may first aim to get basic understanding of a new programming topic by perusing parts of the coursebook, reading and completing examples within the book — all as part of a single coursebook-oriented tactic. To reach the learning goals, students may adjust their tactics over time based on newly set sub-goals, self-evaluations and changing learning context. Therefore, a student’s learning strategy comprises a set of learning tactics and their approach to varying them to reach a larger learning goal. For instance, to complete programming exercises for a week, a student’s strategy may be to first use a coursebook-oriented tactic to get more content knowledge, then upon successful understanding shift to more task-oriented tactic which encompasses completing tasks while using course materials as a reference. Overall, such event-based characterization of learning strategies enables the description of the study process as an enactment of learning actions based on the current conditions. Consequently, this definition allows for the graphical representation of student behavior as sequences of actions, tactics, and strategies over time (e.g., Hadwin et al., 2007).
@@ 2.2. Self-regulated learning in computing education
Silva et al. (2024) also discussed the challenges students face in maintaining motivation and managing time, which reinforces the need for educational interventions to help students better apply SRL strategies. Jovanović et al. (2017) found that students in flipped classrooms often rely on less effective, passive strategies, such as video watching, rather than more active approaches like formative assessments. Similarly, Matcha et al. (2019) and Liao et al. (2019) discussed how different analytic approaches can identify varying learning strategies in online courses, and argued for selecting appropriate methods such as tailoring and scaffolding to better support learners. Similar insights were also discussed by Panadero (2017) who reviewed six SRL models, concluding that these frameworks should be tailored to students’ developmental stages. Focusing their review of SRL models applied in computing education, Loksa et al. (2022) further suggested that researchers should select a framework and concepts that best fit their research context. Greene and Azevedo (2007) discussed the complexity of SRL, advocating for research that considers the interactions between cognitive and task conditions across learning phases.
@@ 2.3. At-risk student detection
More recently, Sghir et al. (2022) noted the increasing use of machine and deep learning models in predicting academic outcomes, paving way for using more performance data in predictions. Waheed et al. (2023) demonstrated high effectiveness of deep learning techniques in predicting students at risk of failing in self-paced online education. Focusing on privacy concerns, Van Petegem et al. (2022) presented a privacy-friendly framework for early detection of students at risk of failing introductory programming courses. Utilizing interpretable machine learning techniques, they achieved high predictive accuracy early in the semester. A further study by Zhidkikh et al. (2024) reproduced these results using the same framework in a different CS1 course, highlighting the generalizability and explainability of at-risk student predictions through submission metadata and noting the importance of student self-evaluations in improving early dropout predictions. Finally, Moreno-Marcos et al. (2020) investigated dropout prediction in self-paced MOOCs incorporating self-regulated learning strategies as predictors. They found that event-based SRL strategies, derived from learners’ interactions with course materials, had high predictive power, outperforming traditional demographic and self-reported measures. Their result emphasizes the importance of considering complex, activity-based features over conventional data in predictive models.
@@
3. Research context and methods
@@ 3.1. Course context
The course followed a weekly schedule, with a suggested learning strategy depicted in Figure 1. Each week began with two lectures for which the students could prepare by reviewing the week’s materials. The lectures were held in-person and online, during which the lecturer and students dynamically interacted through quizzes, interactive examples and discussions on the week’s topic.
Following the lectures, students completed a weekly task set consisting of 10 to 20 graded exercises of varying difficulty levels: introductory tasks (e.g., simulators and quizzes), basic programming tasks, bonus tasks (requiring more work), and guru tasks (requiring seeking extra knowledge beyond the week’s material). Students generally had the freedom to choose which exercises to submit, except for one or two “core tasks” that were mandatory each week. They had access to several learning resources, such as a coursebook, interactive examples, lecture recordings, and documentation. Students were encouraged to use only the course materials and not search the internet, as all necessary and verified materials are available within the course.
@@
In addition, students could attend group study sessions in which they could study and get live help from teaching assistants. During the course, students were also required to participate in three one-on-one supervision sessions with a teaching assistant to assess the progress on the course assignment. Each course week ended with a task review session held after the task deadline, allowing students to compare each other’s answers. After the session, the students were encouraged to review and fix their answers using model answers.

### v2

@@ Abstract.
The introductory programming course (CS1) at the university level is often perceived as particularly challenging, contributing to high dropout rates among Computer Science students. Identifying when and how students encounter difficulties in this course is critical for providing targeted support. This study explores the behavioral patterns of CS1 students at varying dropout risks using self-regulated learning (SRL) as the theoretical framework. Using learning analytics, we analyzed trace logs and task performance data from a virtual learning environment to map resource usage patterns and used student dropout prediction to distinguish between low and high dropout risk behaviors. Data from 47 consenting students were used to carry out the analysis. Additionally, self-report questionnaires from 29 participants enriched the interpretation of observed patterns. The findings reveal distinct weekly learning strategy types and categorize course behavior. Among low dropout risk students, three learning strategies were identified that differed in how students prioritized completing tasks and reading course materials. High dropout risk students exhibited nine different strategies, some representing temporary unsuccessful strategies that can be recovered from, while others indicated behaviors of students on the verge of dropping out. This study highlights the value of combining student behavior profiling with predictive learning analytics to explain dropout predictions and devise targeted interventions. Practical findings of the study can in turn be used to help teachers, teaching assistants and other practitioners to better recognize and address students at the verge of dropping out.
@@ 1. Introduction
Identifying struggling students and those at risk of dropping out is crucial to ensuring these interventions are effective. Academic performance prediction and student dropout detection have been extensively researched (Hellas et al., 2018; Prenkaj et al., 2020). Many studies apply learning analytics techniques to various student data, including demographics (e.g., Naseem et al., 2019), basic activity indicators (e.g., Waheed et al., 2023), outcomes of student assignments (e.g., Zhidkikh et al., 2024b) and self-reports (e.g., Tempelaar et al., 2018a). Notably, Moreno-Marcos et al. (2020) evaluated an effective prediction approach using students’ self-regulation behavior data collected from self-reports and trace logs. Although most student dropout prediction methods demonstrate high accuracy, easy explainability of the predictions remains a challenge (Zhidkikh et al., 2024b).
@@
To increase the effectiveness of the predictions and interventions, it is therefore essential to understand how students study and how they utilize available learning resources. Recent research has increasingly focused on identifying student learning behaviors, emphasizing the self-regulation, metacognition, and motivation aspects (Lorås et al., 2022). For instance, common activity indicators can be used to model students’ learning approaches and to identify student behavior profiles (e.g., Kim et al., 2020; Sharma et al., 2015). Similarly, trace logs from virtual learning environments can reveal patterns in students’ learning resource usage and detect common learning strategies used in a classroom (e.g., Jovanović et al., 2017; Zhidkikh et al., 2023a). Additionally, self-reports, such as questionnaires, learning journals, and interviews may provide insights into students’ use of learning strategies and cognitive processes (e.g., Loksa and Ko, 2016; Loksa et al., 2020a; Silva et al., 2024). However, while these methods help map learning behaviors, they usually do not directly link behavior to early dropout (Zhidkikh et al., 2023a). At most, some studies have compared behaviors of high and low-performing students, often in a post-course setting. For instance, Liao et al. (2019) interviewed 19 CS1 students who took the final CS1 exam and identified general differences in learning paths between higher-performing and lower-performing students.
@@
Motivated by the examples above, this paper aims to deepen the understanding of learning behaviors among high and low-risk students during CS1. The present paper complements prior key studies (Liao et al., 2019; Moreno-Marcos et al., 2020). Specifically, we map students’ learning paths, extending previous work (cf. Liao et al., 2019) by tracking weekly learning behaviors using trace logs collected during the course. Further, we employ dropout prediction to identify at-risk students for every course week (cf. Moreno-Marcos et al., 2020) and use this information to discern weekly learning behavior by the associated dropout risk. To this end, the paper presents a learning analytics-based method to link learning resource usage patterns with dropout predictions, combining tools developed in prior studies (Zhidkikh et al., 2023a, 2024b). A post-course self-report questionnaire was also administered to support the interpretation of the behavioral patterns and to evaluate the potential of the learning analytics used. Both qualitative and quantitative data were inductively explored to gain insights into students’ behaviors at different dropout risk levels. We use self-regulated learning as the foundational theoretical framework widely applied in CSEd research to understand the learning process (Prather et al., 2020). The study ultimately seeks to answer the following research questions:
@@ 2.1. Self-regulated learning and learning strategies
Self-regulation generally refers to the ability to monitor and manage one’s effort, time, motivation and behavior when learning or problem-solving (Pintrich, 2000; Schunk, 2005). While metacognition describes student’s knowledge to self-evaluate learning outcomes and approaches, self-regulation describes the ability to control them (Prather et al., 2020). Therefore, SRL is a strategic learning process in which a student self-regulates by actively setting goals, engaging with the learning resources using various strategies, monitoring progress and refining their approach based on the outcomes and external learning context (Zimmerman, 2008; Winne and Hadwin, 1998; Winne and Marzouk, 2019). Although multiple models detailing the SRL process exist (Panadero, 2017; Loksa et al., 2022b), SRL is often described as a cyclical process consisting of forethought and goal-setting, enacting on the goals and self-controlling, and self-reflection on the outcomes and the process (Zimmerman and Moylan, 2009). While the general SRL process can be modelled, the specific ways a student regulates their behavior can depend highly on the context, such as the subject, student’s prior experiences, the task at hand and even other students (Ben-Eliyahu and Bernacki, 2015).
@@
Various measures of SRL have been applied in academics. Common measures include self-report tools that assess student’s aptitudes for self-regulation (e.g., questionnaires, think-aloud protocols, learning journals), and event measures that map SRL processes via observable, externally collected data (Roth et al., 2016). While some researchers have noted issues related to the reliability of self-reports compared to event-based measures (e.g., Choi et al., 2023; Zimmerman, 2008), a growing number of studies has argued for the combined use of both measures (e.g., Fan et al., 2023b; Lim et al., 2021; Tempelaar et al., 2020b). As Roth et al. (2016) summarize, the triangulation of new and established measures enables researchers to gain new insights into SRL.
A common aspect used to evaluate and study SRL is the use of learning strategies. In a classic conceptualization outlined by Zimmerman and Martinez-Pons (1986), a learning strategy is a set of planned actions, such as rehearsing, seeking assistance, and reviewing course materials, aimed at acquiring information and skills relevant to the goal or task at hand. This definition has been commonly used in SRL research based on self-reports or theoretical frameworks that build on Zimmerman’s work (e.g., Falkner et al., 2014; Silva et al., 2024; Moreno-Marcos et al., 2020; Garcia et al., 2018).
@@
For event-based measures, Winne and Hadwin (1998) proposed a more fine-grained distinction of student actions into learning tactics and learning strategies. The conceptualization was later expanded by Winne (2001) and Winne and Marzouk (2019). In this conceptualization, a learning tactic is a sequence of discrete, atomic actions that a student performs to achieve a sub-goal or a concrete task. A learning strategy, in turn, is a higher-level sequence of tactics that a student employs to reach a larger learning goal and that is guided by the student’s metacognitive monitoring and control. For example, a student may first aim to gain a basic understanding of a new programming topic by perusing parts of the coursebook and reading and completing examples within the book — all as part of a single coursebook-oriented tactic. To complete programming exercises for a week, in contrast, a student’s strategy may be to first use a coursebook-oriented tactic to build content knowledge and then, once that understanding is in place, shift to a more task-oriented tactic that involves completing tasks while using course materials as a reference. To reach the learning goals, students may adjust their tactics over time based on newly set sub-goals, self-evaluations, and changing learning context. In terms of the overall SRL process, Winne (2001) positions learning strategy selection as part of the forethought phase, while learning tactics are dynamically adjusted during the enacting phase in response to the state of the task. Winne and Marzouk (2019) note that tactics are specifically selected to enact on the plans of a learning strategy.
Compared to the classic conceptualization of learning strategies, the tactic–strategy distinction provides a hierarchical view of student actions, where tactics are seen as building blocks of strategies (Winne and Marzouk, 2019). Such an event-based characterization allows describing the study process as an enactment of learning actions based on the current conditions. Consequently, this definition allows the graphical representation of student behavior as sequences of actions, tactics, and strategies over time (Winne and Marzouk, 2019).
The tactic–strategy conceptualization of student behavior has been widely applied in learning analytics research over the past decade. The most prominent operationalization builds on the seminal work of Matcha et al. (2019b), in which learning tactics and strategies are identified using process mining and cluster analysis on student trace logs obtained from an online learning environment. Learning tactics are operationalized as patterns in student action sequences within an uninterrupted period of study, referred to as a learning session. These patterns are extracted from study sessions using cluster analysis on the sequence of actions performed by students. Learning strategies are, in turn, defined as students’ tactic usage patterns across the studied period and are identified through further cluster analysis. This approach has since been replicated and verified in various study contexts, such as flipped classrooms (Uzir et al., 2020), online courses (Fan et al., 2021a), medical education (Piotrkowicz et al., 2021), and secondary education (Zhidkikh et al., 2023a).
While the process mining and cluster analysis approach has been widely adopted, other operationalizations exist. For example, Fincham et al. (2019) analyzed the learning tactics and strategies of first-year engineering students by modeling tactics as a Hidden Markov Model rather than applying cluster analysis to the sequence of actions. Some subsequent studies have also used Hidden Markov Models (Yang et al., 2025; Villalobos et al., 2022), noting their ability to model the sequential nature of learning tactics. Other recent studies have explored network analysis approaches. Fan et al. (2023c) applied ordered network analysis to identify and interpret learning tactics in an online course, with results suggesting that network-based methods may provide deeper characterizations of learning tactics than process mining. Further network-based approaches to tactic–strategy identification include epistemic network analysis (Rakovic et al., 2023) and psychological networks (Saqr et al., 2020).
Overall, there are multiple approaches to implementing event-based measures for SRL when applying the tactic–strategy characterization of Winne and Hadwin (1998). However, it is important to recognize that no single approach yields a definitive description of self-regulated learning in a given context. In a comparative study by Matcha et al. (2019a), process mining, sequence analysis, and network analysis approaches were evaluated for their ability to detect learning tactics and strategies. The study found that while the approaches produced slightly different sets of learning tactics and strategies, the differences were mainly attributable to the varying emphases of the approaches. As such, Matcha et al. (2019a) highlight that what matters more than the exact operationalization of student behavior is how the results are interpreted and grounded in the studied context. In other words, domain knowledge plays a critical role in implementing event-based measures of self-regulation.
@@ 2.2. Self-regulated learning in computing education
Silva et al. (2024) also discussed the challenges students face in maintaining motivation and managing time, which reinforces the need for educational interventions to help students better apply SRL strategies. Jovanović et al. (2017) found that students in flipped classrooms often rely on less effective, passive strategies, such as video watching, rather than more active approaches like formative assessments. Similarly, Matcha et al. (2019a) and Liao et al. (2019) discussed how different analytic approaches can identify varying learning strategies in online courses, and argued for selecting appropriate methods such as tailoring and scaffolding to better support learners. Similar insights were also discussed by Panadero (2017) who reviewed six SRL models, concluding that these frameworks should be tailored to students’ developmental stages. Focusing their review of SRL models applied in computing education, Loksa et al. (2022b) further suggested that researchers should select a framework and concepts that best fit their research context. Greene and Azevedo (2007) discussed the complexity of SRL, advocating for research that considers the interactions between cognitive and task conditions across learning phases.
@@ 2.3. At-risk student detection
More recently, Sghir et al. (2022) noted the increasing use of machine and deep learning models in predicting academic outcomes, paving way for using more performance data in predictions. Waheed et al. (2023) demonstrated high effectiveness of deep learning techniques in predicting students at risk of failing in self-paced online education. Focusing on privacy concerns, Van Petegem et al. (2022) presented a privacy-friendly framework for early detection of students at risk of failing introductory programming courses. Utilizing interpretable machine learning techniques, they achieved high predictive accuracy early in the semester. A further study by Zhidkikh et al. (2024b) reproduced these results using the same framework in a different CS1 course, highlighting the generalizability and explainability of at-risk student predictions through submission metadata and noting the importance of student self-evaluations in improving early dropout predictions. Finally, Moreno-Marcos et al. (2020) investigated dropout prediction in self-paced MOOCs incorporating self-regulated learning strategies as predictors. They found that event-based SRL strategies, derived from learners’ interactions with course materials, had high predictive power, outperforming traditional demographic and self-reported measures. Their result emphasizes the importance of considering complex, activity-based features over conventional data in predictive models.
@@
3. Research context and data
@@ 3.1. Course context
Week Topic Course assignment schedule
1 Introduction to command line, simple C# program
2 Functions, documentation, variables
3 IDE usage, more functions and variables
4 Parameters, functions, introduction to unit tests
5 Conditionals, strings Milestone 1 (Planning)
6 Loops, arrays
7 Multi-dimensional arrays, string splitting
8 Dynamic data structures, intro into object oriented programming Milestone 2 (Prototype)
9 Game development, file handling
10 Recursion, exceptions, function references, lambda expressions
11 Floating point numbers, questions from students Milestone 3 (Final)
Table 1. The syllabus of the studied CS1 course. The course comprises 11 weekly topics that each includes two lectures and exercises. Additionally, the course includes a mini-project assignment that is split into three milestones that students have to present to the teaching assistants.
@@ 11 Floating point numbers, questions from students Milestone 3 (Final)
The course syllabus is outlined in Table 1. The course followed a weekly schedule, with a learning strategy suggested by the course teacher depicted in Figure 1. Each week began with two lectures for which the students were encouraged to prepare by reviewing the week’s materials. The lectures were held in-person and online, during which the lecturer and students interacted through quizzes, interactive examples and discussions on the week’s topic.
Following the lectures, students completed a weekly task set consisting of 10 to 20 graded exercises of varying difficulty levels. All exercises were divided into four main types. Introductory tasks included quizzes and tasks based on visual simulators with the goal to scaffold students into the week’s topic. Basic tasks refer to programming tasks relevant to the week’s topic. For example, basic tasks in week 2 revolved around defining and using functions to draw staircases within the game engine used in the course, and basic tasks in week 6 focused on solving the rainfall problem using arrays and loops. Bonus tasks contained any extra tasks related to the week’s topic and were mainly provided for those interested in additional practice or extra points for a better course grade. For instance, bonus tasks in week 1 included practicing programming with Alice [^1] https://alice.org, a logic quiz and answering the course’s preliminary survey. Finally, guru tasks required students to seek extra knowledge beyond the week’s material and were mainly provided for more skilled students. Guru tasks included, for instance, programming a Conway’s Game of Life simulator and solving Project Euler problems or prior Advent of Code problems. For brevity, we collectively refer to bonus and guru tasks as extra tasks in this paper.
Students generally had the freedom to choose which exercises to submit. However, a small selection of introductory and basic tasks were designated as mandatory core tasks. Submitting answers to the core tasks was a requirement for passing the course to ensure that all students completed at least some tasks from every week. The number of core tasks varied weekly from two to three and they were selected by the course teacher beforehand.
In addition to the weekly tasks, students completed a course assignment throughout the course. The assignment comprised a self-made mini-project, which was usually a game made with the game engine used in the course or a small command-line program. Students were allocated 30 hours of course work for the assignment, which was split into three milestones: a plan document, a working prototype and the finished project. During the course, students were required to participate in three one-on-one supervision sessions with a teaching assistant to assess the progress of the project throughout the milestones (see Table 1). The assignment was evaluated as a whole in the final milestone and was graded as pass/fail. Those whose assignment did not match the passing criteria as evaluated by the TA were required to revise their work and participate in a new one-on-one supervision session.
@@
Throughout the course, the students had access to several learning resources, such as a coursebook, lectures, supplementary examples and ungraded exercises, lecture recordings, and documentation. Students were encouraged to use only the course materials and not search the internet, as all necessary and verified materials are available within the course. In addition, students could attend group study sessions in which they could study and get live help from teaching assistants. Additionally, each course week ended with a task review session held after the task deadline, allowing students to compare each other’s answers. After the session, the students were encouraged to review and fix their answers using model answers provided by the course teacher. Similar to course materials, students were encouraged to attend lectures, group study sessions and task review sessions, but there was no required attendance. In fall 2023, $62.6\%$ of the course students self-reported attending some of the lectures, with an average self-reported attendance of $19.4\%$ per lecture; the lowest reported attendance was $2.8\%$ (week 1, lecture 2) and the highest $38.6\%$ (week 2, lecture 2). In turn, $78.2\%$ of the students were recorded as participating in some of the group study sessions, with an average attendance of $28.9\%$ students per week; the lowest attendance was $14.6\%$ (week 10) and the highest $49.5\%$ (week 7).